If the AM between the roots of a quadratic equation is 8 and the GM is 5, then the equation is

Grade:12

1 Answers

bharat bajaj IIT Delhi
9 years ago
AM = 8 ⇒ Sum of roots = 2 × 8 = 16
GM = 5 ⇒ Product of roots = 52= 25
⇒ x2- 16x + 25 = 0
Thanks & Regards
Bharat Bajaj
askiitian faculty
IIT Delhi